You are here: Foswiki>Commissioning Web>Com20100703 (03 Jul 2010, KatherineRueff)Edit Attach

Commissioning: 3 July 2010 UT Night Lost due to weather conditions.

Observer: Pascale H (LBTO) Doug M (Tucson)
Telescope Operator: Steve A (LBTO)
Telescope Support: (LBTO)
Instrument Support: Tim S (LBTO)
Adaptive Secondary Operator: Steve A (LBTO) Marco X (Firenze)


Started taking pointing model data until weather became too cloudy and started raining, we were never able to open again due to rain and cloud cover.


Waiting for clouds and rain to dissipate.


04:41 Open. Humidity= 66.3%, Air Temp= 13.3 C, Wind Speed= 2.0 m/s
04:47 AO system Powered off without reason, hexapod wasn't responding, had to reinitialize it.
04:51 Turned on power again, Loaded program, and Pascale set AO shell.


04:56 pointto ACT0233 Tel="ACQUIRE" gs=0 PA=0 PAM-
04:57 preset failed due to register is still set to LUCIFER, changed back to IRTC
04:48 preset failed due to star below horizon.
04:59 pointto ACT0232 Tel="ACQUIRE" gs=0 PA=0 PAM-, EL = 56
05:00 Error in preset:
Resmsg[0] = PresetTelescope result status: Error
Resmsg[1] = [2]
Resmsg[2] = dx.terc: [adjustCurrFocalStation] Not at a focal station
05:01 Steve calls Paul to chat about focal station OSS problem.
05:05 Resent preset, failed again, Steve works with Paul to fix problem.
05:10 Problem fixed by Paul setting OSS to yesterday's version since he updated it today, resent preset.
05:15 Steve fixes coma and astigmatism, the coma is very bad.
05:19 IE=6, CA= -4
05:22 pointto BS9177 Tel="ACTIVE" gs=0; Waited for collimation, looking through clouds, so changing star
05:28 pointto BS9164 Tel="ACTIVE" gs=0
05:31 Taking a long time to collimate, scope is jumping from seeing 0.8" to 1/3"

Pointing Model

05:40 Turned on pointing log in PCSGUI. File: /home/telescope/20100703.ptlog 05:41 cp ~LBTO/POINTING/* . to copy program files to current directory. Ran prepdir
05:42 epar gopoint, changed file to gopointjuly032010, agwNum= 2, exptime= 0.1sec, Nimage= 100, and ran script.
05:44 WT10_380 target is saturated and has astigmatism of 3"
05:46 Changing exp time to 0.01 sec.
05:49 Rerun gopoint, star is still too bright, full saturating WFS.
05:50 Changed exp time to 0.0001, rerun script.
05:52 Image still saturated, but letting script run still
  • irtc.20100703.055250.fits with cube 2 (This image is empty of a star)
05:55 Getting error:
Resmsg[0] = PresetTelescope result status: Error
Resmsg[1] = [22]
Resmsg[2] = couldn't adjust exposure time to meet low counts requirement. star too dim or guide rate too high.
Resmsg[3] = maximum exposure time is 2100 ms at the given guide rate, estimated counts are 260
Resmsg[4] = adjusted exposure time to 2100 ms
Resmsg[5] = guiding was not stable within 6 cycles to examine the log: page LOGs/pointto_055304.log
05: 56 stopped script


06:05 Closed due to cloud cover. Called Doug to discuss problems with pointing model.
06:47 After discussion with Doug we are changing the epar for gopoint to, exp time= 0.1, active= no.

11:18 Steve officially closes for the night.

-- DougMiller - 24 Jun 2010
Topic revision: r4 - 03 Jul 2010, KatherineRueff
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback